OpenStudy (anonymous):

can some1 plz help me with standard notation

OpenStudy (anonymous):

sure, what is it ?

OpenStudy (anonymous):

5 to the 5th power

OpenStudy (anonymous):

\[5^{5} = 5*5*5*5*5 =125\]

OpenStudy (anonymous):

\[5 \times 5 \times 5 \times 5 \times 5 \times 5 = 5^5\]

OpenStudy (anonymous):

? im in 5th grade

OpenStudy (anonymous):

alright :) and after u get done i will show u me and MY sister :)

OpenStudy (anonymous):

she drool A LOT!!

OpenStudy (anonymous):

awwwwwwwwwwwww!!!!!!! she is adorable!!! though there is no such thing as a ugly baby its just not possible :)

OpenStudy (anonymous):

lolz wanna see my other lil sis

OpenStudy (anonymous):

totally!! :) hold on i cant find the pics of me and my sister >.< LOL

OpenStudy (anonymous):

me and my sis im the blonde ^_^

OpenStudy (anonymous):

aaaaaawwwww

OpenStudy (anonymous):

^_^ not my best picture :P but ya lol

OpenStudy (anonymous):

lolz

OpenStudy (anonymous):

im bored and dont feel like doin math

OpenStudy (anonymous):

lol ur sisters are sooooo cute!!!! i bet ur gorgeous!!!

OpenStudy (anonymous):

omg i am!!! wanna see me!

OpenStudy (anonymous):

haha!! okie dokie sure!! :)

OpenStudy (anonymous):

omg i cnt find 1 !!!

OpenStudy (anonymous):

:) u wanna see me? :)

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

okie dokie hold on a second... >.< finding a not horrible picture of me lol

OpenStudy (anonymous):

urghhhh >.<

OpenStudy (anonymous):

im not so gorgeous >.<

OpenStudy (anonymous):

omkg ur pretty!!!

OpenStudy (anonymous):

ok brb i gotta take 1 lolz

OpenStudy (anonymous):

ya watever dont lie :)

OpenStudy (anonymous):

no u are!!!

OpenStudy (anonymous):

lmao.. why are yu two chatting on a question ? lol

OpenStudy (anonymous):

idk lolz

OpenStudy (anonymous):

oh no non no

OpenStudy (anonymous):

srry i was sleeping

OpenStudy (anonymous):

:) lol ur very cute girl!!! :)

OpenStudy (anonymous):

and math guy we are chatting on here to share our pics :)

OpenStudy (anonymous):

why dont u share urs!! :D lmao im jk :)

OpenStudy (anonymous):

haha i see.. and nah thakns (;

OpenStudy (anonymous):

too bad >.< im jk :) lol i already have somone anyways :)

OpenStudy (anonymous):

the babies are soooooooo ca-ute :) lol i was totally freaking out that dude!!! LMAO

OpenStudy (anonymous):

des email me